OpenStudy (anonymous):

HEY GUYS WHAT'S UP? I have a question and I need it explained step by step for my test, PLEASE HELP! The question is factor out the greatest common factor from the expression: 20x^8+30x^7-15x^4?

jimthompson5910 (jim_thompson5910):

The GCF of 20, 30, and -15 is 5 (this is the largest number that can go into all 3) the GCF of x^8, x^7, and x^4 is x^4 (same reason) So the GCF of 20x^8 , 30x^7 and -15x^4 is 5x^4 Factor out 5x^4 by dividing each term by this GCF to get 5x^4 ( 4x^4 + 6x^3 - 3) Let me know if you need me to explain how I got the stuff inside the parenthesis

OpenStudy (anonymous):

Yes I do please Jim just one more time

jimthompson5910 (jim_thompson5910):

First term: (20x^8)/(5x^4) = (20/5)x^(8-4) = 4x^4 Second term: (30x^7)/(5x^4) = (30/5)x^(7-4) = 6x^3 Third term: (-15x^4)/(5x^4) = (-15/5)x^(4-4) = -3

jimthompson5910 (jim_thompson5910):

each time, I'm dividing an original term by the GCF to get the new term

OpenStudy (anonymous):

okay I need to ask you a question

jimthompson5910 (jim_thompson5910):

what's that

OpenStudy (anonymous):

when I am dividing, do I use the multiplication rules to get the positive or negatives?

OpenStudy (anonymous):

and negatives

jimthompson5910 (jim_thompson5910):

yes, something like -15/5 = -3 you divide 15 by 5 to get 3, then you make it negative since one number is negative

jimthompson5910 (jim_thompson5910):

is that what you're asking about?

OpenStudy (anonymous):

so its just like multiplying the signs, yes thats what I'm asking

jimthompson5910 (jim_thompson5910):

yes, dividing uses the same rules as multiplying when it comes to positives and negatives

OpenStudy (anonymous):

AWESOME

jimthompson5910 (jim_thompson5910):

(positive)/(positive) = positive (negative)/(negative) = positive (negative)/(positive) = negative (positive)/(negative) = negative

OpenStudy (anonymous):

THANK YOU JIM!!! YOUR THE BEST!!!:)

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

hey buddy, I have a question, I'm not sure if I got it correct or not

OpenStudy (anonymous):

are you still here??

jimthompson5910 (jim_thompson5910):

what's the question